Problem 5-35 (LO 5-2) On July 1 of year 1, Riverside Corp. (RC), a calendar-year taxpayer, acquired the assets of another busReq A Req B What book-tax difference associated with its goodwill should RC report in year 1? Is it favorable or unfavorable?Req A Req B What book-tax difference associated with its goodwill should RC report in year 2? Is it favorable or unfavorable?
Answer #1

Solution:

a:

Book-tax difference $ 40,000
Is it favorable or unfavorable? Favorable
Is it permanent or temporary? Temporary

According to Tax authorities  RC should amortize $1,200,000 of goodwill over a period of 15 years. The depreciation method to be used will be straight-line method.

Amortization amount for year 1 will be for half the year i.e. 6 months. Now 15 years = 15 x 12 = 180 months.

Amount = (1,200,000/180) x 6 = $ 40,000. This amount is for tax purposes.

As determined by RC there is no deduction for book purposes . Thus the difference is 40,000 - 0 = $40,000. This is a favorable difference as no deduction has been made for book purpose and also it is temporary.

b.)

Book-tax difference $ 120,000
Is it favorable or unfavorable? Unfavorable
Is it permanent or temporary? Temporary

For tax purposes,  Amortization amount for year 2 = 1,200,000/15 years = $80,000.

Amount for book purpose = 200,000 (given)

Difference = 200,000 - 80,000 = 120,000

This amount is unfavorable as amortization for book purpose is greater than the amount for tax purposes. The difference is temporary.